# Amortized Inference of Causal Models via Conditional Fixed-Point Iterations

### v3

@@ E.1 Experiments with $n_{\mathcal{D_{\mathrm{test}}}}=100$
In this section we benchmark Cond-FiP against the baselines for the scenario when test datasets in the input context have smaller sample size ($n_{\mathcal{D_{\mathrm{test}}}}=100$) as compared to the train datasets ($n_{\mathcal{D_{\mathrm{test}}}}=400$) in Appendix C.
We report the results for the task of noise prediction, sample generation, and interventional generation in Table 10Table 11, and Table 12 respectively. We find that Cond-FiP exhibits superior generalization as compared to baselines. For example, in the case of RFF IN, Cond-FiP is even better than FiP for all the tasks! This can be attributed to the advantage of amortized inference; as the sample size in test dataset decreases, the generalization of baselines would be affected a lot since they require training from scratch on these datasets. However, amortized inference methods would be impacted less as they do not have to trained from scratch, and the inductive bias learned by them can help them generalize even with smaller input context.
@@
Cond-FiP 100 $0.07$ $(0.0)$ $0.13$ $(0.01)$ $0.08$ $(0.01)$ $0.17$ $(0.01)$
Table 10: Results for Noise Prediction with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=100$). We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Each test dataset contains $100$ samples, as opposed to $400$ samples in Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@
Cond-FiP 100 $0.11$ $(0.01)$ $0.35$ $(0.07)$ $0.14$ $(0.02)$ $0.63$ $(0.07)$
Table 11: Results for Sample Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=100$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Each test dataset contains $100$ samples, as opposed to $400$ samples in Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@
Cond-FiP 100 $0.12$ $(0.02)$ $0.32$ $(0.07)$ $0.13$ $(0.01)$ $0.58$ $(0.07)$
Table 12: Results for Interventional Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=100$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variable. Each test dataset contains $100$ samples, as opposed to $400$ samples in Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@ E.2 Experiments with $n_{\mathcal{D_{\mathrm{test}}}}=50$
We conduct more experiments for the smaller sample size scenarios, where decrease the sample size even further to $n_{\mathcal{D_{\mathrm{test}}}}=50$ samples. We report the results for the task of noise prediction, sample generation, and interventional generation in Table 13Table 14, and Table 15 respectively. We find that baselines perform much worse than Cond-FiP for the all different SCM distributions, highlighting the efficacy of Cond-FiP for inferring causal mechanisms when the input context has smaller sample size. Note that there were issues with training DoWhy for such a small dataset, hence we do not consider them for this scenario.
@@
Cond-FiP 100 $0.09$ $(0.01)$ $0.16$ $(0.01)$ $0.09$ $(0.01)$ $0.18$ $(0.01)$
Table 13: Results for Noise Prediction with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=50$). We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Each test dataset contains $50$ samples, as opposed to $400$ samples in Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@
Cond-FiP 100 $0.16$ $(0.01)$ $0.42$ $(0.07)$ $0.22$ $(0.01)$ $0.65$ $(0.06)$
Table 14: Results for Sample Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=50$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Each test dataset contains $50$ samples, as opposed to $400$ samples in Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@
Cond-FiP 100 $0.12$ $(0.01)$ $0.33$ $(0.07)$ $0.14$ $(0.01)$ $0.58$ $(0.07)$
Table 15: Results for Interventional Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=50$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variable. Each test dataset contains $50$ samples, as opposed to $400$ samples in Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@ Appendix F Experiments without True Causal Graph on AVICI Benchmark
Results in Appendix C (Table 1Table 2Table 3) require the knowledge of true graph (${\bm{\mathcal{G}}}$) as part of the input context to Cond-FiP. In this section we conduct where we don’t provide the true graph in the input context, rather we infer the graph $\hat{\bm{\mathcal{G}}}$ using an amortized causal discovery approach (AVICI [Lorch et al., 2022]) from the observational data $D_{\bm{X}}$. We chose AVICI for this task since it can enable to amortized inference of causal graphs, hence allowing the combined pipeline of AVICI + Cond-FiP can perform amortized inference of SCMs. More precisely, AVICI infers the graph from a novel instance ${\bm{\mathcal{G}}}$ from input context $D_{\bm{X}}$ without updating any parameters, and we pass $(\hat{\bm{\mathcal{G}}},D_{\bm{X}})$ as the input context for Cond-FiP. Therefore, for any $\bm{z}\in\mathbb{R}^{d}$, Cond-FiP ( ${\mathcal{T}}(\bm{z},D_{\bm{X}},\hat{\bm{\mathcal{G}}})$) aims to replicate the functional mechanism ${\bm{F}}({\bm{z}})$ of the underlying SCM.
The results for benchmarking Cond-FiP with inferred graphs using AVICI for the task of noise prediction, sample generation, and interventional generation are provided in Table 16Table 17, and Table 18 respectively. For a fair comparison, the baselines FiP, DECI, and DoWhy also use the inferred graph ($\hat{\bm{\mathcal{G}}}$) by AVICI instead of the true graph (${\bm{\mathcal{G}}}$). We find that Cond-FiP remains competitive to baselines even for the scenario of unknown true causal graph. Hence, our training procedure can be extended for amortized inference of both causal graphs and causal mechanisms of the SCM.
@@
Cond-FiP 100 $0.48$ $(0.06)$ $0.34$ $(0.03)$ $0.64$ $(0.03)$ $0.49$ $(0.04)$
Table 16: Results for Noise Prediction without True Graph. We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Unlike experiments in Table 1, the true graph ${\bm{\mathcal{G}}}$ is not present in input context, rather its inferred via AVICI [Lorch et al., 2022]. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph.
@@
Cond-FiP 100 $0.64$ $(0.06)$ $0.49$ $(0.06)$ $0.68$ $(0.02)$ $0.63$ $(0.05)$
Table 17: Results for Sample Generation without True Graph. We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Unlike experiments in Table 2, the true graph ${\bm{\mathcal{G}}}$ is not present in input context, rather its inferred via AVICI [Lorch et al., 2022].. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph.
@@
Cond-FiP 100 $0.64$ $(0.06)$ $0.49$ $(0.06)$ $0.7$ $(0.04)$ $0.62$ $(0.05)$
Table 18: Results for Interventional Generation without True Graph. We compare Cond-FiP against the baselines for the task of interventional data from the input noise variable. Unlike experiments in Table 3, the true graph ${\bm{\mathcal{G}}}$ is not present in input context, rather its inferred via AVICI [Lorch et al., 2022]. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph.
@@ Appendix G Ablation Study on AVICI benchmark
G.1 Ablation Study of Encoder
@@
Note that for the training the subsequent decoder, we sample SCMs with both linear and rff causal mechanisms as in the main results ( Table 1Table 2, and Table 3). Note that in the main results, the encoder was trained by sampling SCMs with both linear and rff functional relationships. Hence, this ablation helps us to understand whether the strategy of training encoder on mixed functional relationships can bring more generalization to the amortization process, or if we should have trained encoders specialized for linear and non-linear functional relationships.
We present our results of the ablation study for the task of noise prediction, sample generation, and interventional generation in Table 19Table 20Table 21 respectively. Our findings indicate that Cond-FiP is robust to the choice of encoder training strategy! Even though the encoder for Cond-FiP (RFF) was only trained on data from non-linear SCMs, its generalization performance is similar to Cond-FiP where the encoder was trained on data from both linear and non-linear SCMs.
@@
G.2 Ablation Study of Decoder
@@
Note that in the main results (Table 2Table 3) we show the performances of Cond-FiP trained by sampling SCMs with both linear and non-linear causal mechanisms. Hence, this ablations helps us to understand whether the strategy of training on mixed causal mechanisms can bring more generalization to the amortization process, or if we should have trained decoders specialized for linear and non-linear functional relationships.
@@
Appendix H Experiments on CSuite with Complex Noise Distributions
(a) Noise Prediction
(b) Sample Generation
(c) Interventional Generation
Figure 7: We compare Cond-FiP against the baselines for the different evaluation tasks on the Large Backdoor and Weak Arrow datasets from the CSuite benchmark, where the noise distribution is modified to be a multi-modal gaussian mixture model. We experiment with $6$ different cases of the noise distribution for each dataset. The y-axis denotes the RMSE for the respective tasks across the $12$ scenarios (datasets & noise distribution). Results indicate that Cond-FiP can generalize to instances with more complex noise distributions like gaussian mixture models.
  • Noise is sampled with equal probability from either $N(-2,1)$ and $N(2,1)$.
  • Noise is sampled with equal probability from either $N(-2,2)$ and $N(2,2)$.
  • Noise is sampled with equal probability from either $N(-2,1)$ and $N(2,2)$.
  • Noise is sampled with equal probability from either $N(-5,1)$ and $N(5,1)$.
  • Noise is sampled with equal probability from either $N(-5,2)$ and $N(5,2)$.
  • Noise is sampled with equal probability from either $N(-5,1)$ and $N(5,2)$.
This leads to a total of $12$ experimental setting with $6$ different GMM noise distribution for both the Large Backdoor and Weak Arrow datasets from the CSuite benchmark. Results in Figure 7 demonstrate that Cond-FiP remains competitive with baselines across all tasks. Importantly, while baselines were trained from scratch for each specific gaussian mixture noise distribution, Cond-FiP was pretrained only on gaussian noise and generalizes effectively to settings with GMM noise distribtion.
Appendix I Experiments on Real World Benchmark
DoWhy $0.015$ $0.014$ $0.005$
DECI $0.014$ $0.005$ $0.005$
FiP $0.015$ $0.005$ $0.005$
Cond-FiP $0.013$ $0.005$ $0.005$
Table 24: Results for Sachs dataset. We benchmark Cond-FiP against the baselines for the task of generating observational data on the real world Sachs benchmark. Each cell reports the MMD, and we also report the reconstruction error for all of the methods. Results indicate that Cond-FiP matches the performance of baselines trained from scratch.
We use the real world flow cytometry dataset [Sachs et al., 2005] to benchmark Cond-FiP againts the baselines. This dataset contains $n\simeq 800$ observational samples expressed in a $d=11$ dimensional space, and the reference (true) causal graph. We split this into context $D^{\text{context}}_{\bm{X}}\in{\mathbb{R}}^{n_{\text{context}}\times d}$ and queries $D^{\text{query}}_{\bm{X}}\in{\mathbb{R}}^{n_{\text{query}}\times d}$, each of size $n_{\text{context}}=n_{\text{query}}=400$. Note that the context dataset is to used to train the baselines and obtain dataset embedding for Cond-FiP, while the query dataset is used for evaluation of all the methods.
Since we don’t have access to the true causal mechanisms, we cannot compute RMSE for noise prediction or sample generation like we did in our experiments with synthetic benchmarks. Instead for each method, we obtain the noise predictions $\widehat{D^{\text{context}}_{\bm{N}}}$ on the context, and use it to fit a gaussian distribution for each component (node). Then we use the learned gaussian distribution to sample new noise variables, $\widehat{D^{\text{query}}_{\bm{N}}}$, which are mapped to the observations as per the causal mechanisms learned by each method, $\widehat{D^{\text{query}}_{\bm{X}}}$. Finally, we compute the maximum mean discrepancy (MMD) distance between $\widehat{D^{\text{query}}_{\bm{X}}}$ and $D^{\text{query}}_{\bm{X}}$ as metric to determine whether the method has captured the true causal mechanisms. For consistency, we also evaluate the reconstruction performances of the models by using directly the inferred noise from context $\widehat{D^{\text{context}}_{\bm{N}}}$ from the models, and then compute MMD between their reconstructed data ($\widehat{D^{\text{context}}_{\bm{X}}}$) and the query data ($D^{\text{query}}_{\bm{X}}$).
Table 24 presents our results, where for reference we also report the MMD distance between samples from the context and query split, which should serve as the gold standard since both the datasets are sampled from the same distribution. We find that Cond-FiP is competitive with the baselines that were trained from scratch. Except DoWhy, the MMD distance with reconstructed samples from the methods are close to oracle performance.
No Interventional Generation Results. Note that Cond-FiP (and the other baselines considered in this work) only supports hard interventions while the interventional data available for Sachs are soft interventions (i.e. the interventional operations applied are unknown). Hence, we are unable to provide a comprehensive evaluation of Cond-FiP (as well as the other baselines) for interventional predictions on Sachs.
Appendix J Comparing Cond-FiP with CausalNF
We also compare Cond-FiP with CausalNF [Javaloy et al., 2023] for the task of noise prediction (Table 25) and sample generation (Table 26). The test datasets consist of $n_{\text{test}}=400$ samples, exact same setup as in our main results (Table 1, Table 2, and Table 3). To ensure a fair comparison, we provided CausalNF with the true causal graph.

### v4

@@ E.1 Experiments with $n_{\mathcal{D_{\mathrm{test}}}}=100$
In this section we benchmark Cond-FiP against the baselines for the scenario when test datasets in the input context have smaller sample size ($n_{\mathcal{D_{\mathrm{test}}}}=100$) as compared to the train datasets ($n_{\mathcal{D_{\mathrm{test}}}}=400$) in Appendix˜C.
We report the results for the task of noise prediction, sample generation, and interventional generation in Table˜12Table˜13, and Table˜14 respectively. We find that Cond-FiP exhibits superior generalization as compared to baselines. For example, in the case of RFF IN, Cond-FiP is even better than FiP for all the tasks! This can be attributed to the advantage of amortized inference; as the sample size in test dataset decreases, the generalization of baselines would be affected a lot since they require training from scratch on these datasets. However, amortized inference methods would be impacted less as they do not have to trained from scratch, and the inductive bias learned by them can help them generalize even with smaller input context.
@@
[…]
Table 12: Results for Noise Prediction with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=100$). We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Each test dataset contains $100$ samples, as opposed to $400$ samples in Table˜3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@
[…]
Table 13: Results for Sample Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=100$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Each test dataset contains $100$ samples, as opposed to $400$ samples in Table˜4. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@
[…]
Table 14: Results for Interventional Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=100$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variable. Each test dataset contains $100$ samples, as opposed to $400$ samples in Table˜5. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@ E.2 Experiments with $n_{\mathcal{D_{\mathrm{test}}}}=50$
We conduct more experiments for the smaller sample size scenarios, where decrease the sample size even further to $n_{\mathcal{D_{\mathrm{test}}}}=50$ samples. We report the results for the task of noise prediction, sample generation, and interventional generation in Table˜15Table˜16, and Table˜17 respectively. We find that baselines perform much worse than Cond-FiP for the all different SCM distributions, highlighting the efficacy of Cond-FiP for inferring causal mechanisms when the input context has smaller sample size. Note that there were issues with training DoWhy for such a small dataset, hence we do not consider them for this scenario.
@@
[…]
Table 15: Results for Noise Prediction with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=50$). We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Each test dataset contains $50$ samples, as opposed to $400$ samples in Table˜3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@
[…]
Table 16: Results for Sample Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=50$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Each test dataset contains $50$ samples, as opposed to $400$ samples in Table˜4. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@
[…]
Table 17: Results for Interventional Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=50$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variable. Each test dataset contains $50$ samples, as opposed to $400$ samples in Table˜5. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
@@ Appendix F Experiments without True Causal Graph on AVICI Benchmark
[…]
The results for benchmarking Cond-FiP with inferred graphs using AVICI for the task of noise prediction, sample generation, and interventional generation are provided in Table˜18Table˜19, and Table˜20 respectively. For a fair comparison, the baselines FiP, DECI, and DoWhy also use the inferred graph ($\hat{\bm{\mathcal{G}}}$) by AVICI instead of the true graph ($\bm{\mathcal{G}}$}). We find that Cond-FiP remains competitive to baselines even for the scenario of unknown true causal graph. Hence, our training procedure can be extended for amortized inference of both causal graphs and causal mechanisms of the SCM.
@@
[…]
Table 18: Results for Noise Prediction without True Graph. We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Unlike experiments in Table˜3, the true graph $\bm{\mathcal{G}}$} is not present in input context, rather its inferred via AVICI (Lorch et al., 2022). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph.
@@
[…]
Table 19: Results for Sample Generation without True Graph. We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Unlike experiments in Table˜4, the true graph $\bm{\mathcal{G}}$} is not present in input context, rather its inferred via AVICI (Lorch et al., 2022).. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph.
@@
[…]
Table 20: Results for Interventional Generation without True Graph. We compare Cond-FiP against the baselines for the task of interventional data from the input noise variable. Unlike experiments in Table˜5, the true graph $\bm{\mathcal{G}}$} is not present in input context, rather its inferred via AVICI (Lorch et al., 2022). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph.
F.1 Analyzing Cond-FiP’s Sensitivity to Causal Graph Structure
To better understand the ability of Cond-FiP to capture graph-specific structure, we conduct a systematic analysis by introducing random perturbations ($p$) to the true causal graph. Specifically, we randomly remove a proportion $p$ of the true edges, such that on average $p\times\text{total edges}$ are missing. We report results for sample generation in the challenging OOD setting (dd=100 & RFF OUT).
Our findings (Table 21) show that Cond-FiP’s performance shows significant difference after we remove 5 edges $(p=0.05)$. Across all tested levels of perturbation, Cond-FiP is competitive with FiP, a baseline trained from scratch for each scenario. These results demonstrate that Cond-FiP exhibits robustness to moderate errors in the input causal graph, comparable to a baseline trained from scratch. This provides more evidence to our claim that Cond-FiP learns causal mechanisms during training and can adapt to new contexts at test time by inferring functions that best explain the available information in the context (input graph and observations), even when the input causal graph is inaccurate.
$p=0$ $p=0.01$ $p=0.02$ $p=0.05$ $p=0.1$
FiP $0.55$ $(0.08)$ $0.55$ $(0.08)$ $0.57$ $(0.08)$ $0.62$ $(0.08)$ $0.68$ $(0.08)$
[…]
Table 21: Robustness to causal graph errors. We compare Cond-FiP for varying levels of corruption ($p$) to the true causal graph by randomly removing a proportion $p$ of edges. Results show that Cond-FiP remains competitive with FiP, even as the input graph becomes increasingly inaccurate, demonstrating robustness to moderate structural errors and its ability to utilize the given imperfect causal information.
@@ Appendix G Ablation Study on AVICI benchmark
G.1 Analyzing the Effect of Pretraining Scale
To better understand the scaling of Cond-FiP w.r.t the pretraining data, we conduct experiments at smaller scales, with a total of $1e5,4e5,1e6$ SCMs, as opposed to using $4e6$ SCMs in our main body results.
Table 22 reports sample-generation accuracy for the $d=20$ setting. As expected, Cond-FiP benefits consistently from additional pretraining data: all four metrics (LIN IN, RFF IN, LIN OUT, RFF OUT) improve monotonically with scale. The most pronounced gains appear when increasing the pretraining size from $1e5$ to $4e5$, while improvements become more incremental beyond $1e6$. These results highlight that Cond-FiP continues to leverage larger synthetic datasets, though the returns gradually diminish as scale increases.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
Cond-FiP($1e5$) 20 $0.11$ $(0.01)$ $0.32$ $(0.06)$ $0.16$ $(0.03)$ $0.44$ $(0.04)$
Cond-FiP($4e5$) 20 $0.07$ $(0.01)$ $0.28$ $(0.06)$ $0.11$ $(0.01)$ $0.37$ $(0.03)$
Cond-FiP($1e6$) 20 $0.07$ $(0.01)$ $0.25$ $(0.06)$ $0.09$ $(0.01)$ $0.33$ $(0.03)$
Cond-FiP($4e6$) 20 $0.05$ $(0.01)$ $0.24$ $(0.06)$ $0.07$ $(0.01)$ $0.30$ $(0.03)$
Table 22: Effect of Pretraining Scale. Sample generation performance of Cond-FiP ($d=20$) under varying amounts of pretraining data. Larger synthetic SCM pretraining datasets lead to consistent improvements across all scenarios, with diminishing but still positive gains beyond $1e6$ SCMs.
G.2 Ablation Study of Encoder
@@
Note that for the training the subsequent decoder, we sample SCMs with both linear and rff causal mechanisms as in the main results ( Table˜3Table˜4, and Table˜5). Note that in the main results, the encoder was trained by sampling SCMs with both linear and rff functional relationships. Hence, this ablation helps us to understand whether the strategy of training encoder on mixed functional relationships can bring more generalization to the amortization process, or if we should have trained encoders specialized for linear and non-linear functional relationships.
We present our results of the ablation study for the task of noise prediction, sample generation, and interventional generation in Table˜23Table˜24Table˜25 respectively. Our findings indicate that Cond-FiP is robust to the choice of encoder training strategy! Even though the encoder for Cond-FiP (RFF) was only trained on data from non-linear SCMs, its generalization performance is similar to Cond-FiP where the encoder was trained on data from both linear and non-linear SCMs.
@@
G.3 Ablation Study of Decoder
@@
Note that in the main results (Table 4Table˜5) we show the performances of Cond-FiP trained by sampling SCMs with both linear and non-linear causal mechanisms. Hence, this ablations helps us to understand whether the strategy of training on mixed causal mechanisms can bring more generalization to the amortization process, or if we should have trained decoders specialized for linear and non-linear functional relationships.
@@
Appendix H Comparing Cond-FiP with CausalNF
We also compare Cond-FiP with CausalNF (Javaloy et al., 2023) for the task of noise prediction (Table˜28) and sample generation (Table˜29). The test datasets consist of $n_{\text{test}}=400$ samples, exact same setup as in our main results (Table 3, Table 4, and Table 5). To ensure a fair comparison, we provided CausalNF with the true causal graph.